# Mamba in Speech: Towards an Alternative to Self-Attention

### v4

@@ D.2 Ablation study of FFN and Residual for ExtBiMamba
D.3 InnBiMamba vs. ExtBiMamba
In Tables 8 and 17, we present the performance of InnBiMamba and ExtBiMamba in Transformer and Conformer frameworks. We found that ExtBiMamba consistently outperforms InnBiMamba across all datasets. Given that InnBiMamba typically has fewer parameters than ExtBiMamba, we increase the number of layers with InnBiMamba to match ExtBiMamba for fair comparison. Again, ExtBiMamba performs better than InnBiMamba.
| Method | Params (M) | dev | test | Params (M) | man | sge |
| Transformer | 30.86 | 13.7 | 13.1 | 29.86 | 17.7 | 24.5 |
| MHSA InnBiMamba | 34.80 | 20.2 | 19.5 | 33.81 | 18.4 | 26.0 |
| MHSA ExtBiMamba | 44.24 | 18.7 | 18 | 43.25 | 17.2 | 24.3 |
[…]
| MHSA InnBiMamba | 50.10 | 12.7 | 12.2 | 49.11 | 17.1 | 23.8 |
| MHSA ExtBiMamba | 56.40 | 12.3 | 11.5 | 55.40 | 16.6 | 23.4 |
| Conformer | ConInnBiMamba | ConExtBiMamba | |
| mins/epoch | 21.4 | 18.3 | 19.8 |
| RTF | 0.179 | 0.174 | 0.177 |
Table 18 compares the training and inference speeds of Conformer, ConInnBiMamba, and ConExtBiMamba. The results show that ConInnBiMamba and ConExtBiMamba train and infer faster than the Conformer, confirming the training and inference efficiency of the ConBiMamba architectures. Among them, ConInnBiMamba is slightly faster than ConExtBiMamba.
Appendix E Ablation Study on ConExtBiMamba for ASR Task
E.1 Effect of Hyper-Parameter for BiMamba in ConExtBiMamba
The default hyper-parameters for ConExtBiMamba are defined as: the SSM state dimension , the state expansion factor , and the local convolution width . From Table 19, we find that increasing the state expansion factor or decreasing the local convolution width has little impact on the performance of ConExtBiMamba.
We further investigate the impact of the matrix on performance. The matrix plays a crucial role in both S4 and Mamba. In the original Mamba paper, it is initialized using a real-valued diagonal matrix. However, ensuring randomness is important in deep learning [58]. To explore this, we examine three types of matrices: a real-valued diagonal matrix, a completely randomized matrix, and a real-valued diagonal matrix initialized by multiplying each element with noise generated from a Gaussian distribution. Through our experiments, we discovered that the diagonal matrix with Gaussian noise yielded the best results.
| Method | dev (↓) | dev other (↓) | test (↓) | test other (↓) |
| ConvExtBiMamba with Gaussian Noise | 5.9 | 17.1 | 6.0 | 17.3 |
| 5.9 | 17.1 | 6.0 | 17.3 | |
| 6.0 | 17.1 | 6.2 | 17.3 | |
| 5.9 | 17.1 | 6.0 | 17.3 | |
| Random Noise A matrix | 6.0 | 17.2 | 6.2 | 17.4 |
| default A matrix | 6.1 | 17.2 | 6.2 | 17.5 |
E.2 Effect of Hyper-Parameters for ConExtBiMamba
We conduct ablation experiments on Swish Activation, Macaron-style feed-forward layers, positional encoding, and dropout. From Table 20, it is apparent that Swish Activation and Macaron-style feed-forward layers enhance the model’s performance, while positional encoding and dropout do not have a significant impact. We suggest that the reason positional encoding, a crucial component in Transformer-based models, does not significantly affect the ConExtBiMamba model is twofold. Firstly, Mamba independently models each input, which suggests that the model may better differentiate between different positions. Secondly, because Mamba operates similarly to an RNN, the dependencies between inputs inherently embed some positional information into the model, thus diminishing the impact of positional encoding.
| Method | dev (↓) | dev other (↓) | test (↓) | test other (↓) |
| ConExtBiMamba | 5.9 | 17.1 | 6.0 | 17.3 |
| Macaron | 6.1 | 17.7 | 6.3 | 17.9 |
| Swish | 6.2 | 17.7 | 6.4 | 18.1 |
| Dropout | 6.0 | 17.1 | 6.0 | 17.4 |
| Positional Encoding | 6.0 | 17.1 | 6.0 | 17.3 |
E.3 Performance on Extremely Small Datasets
To evaluate model robustness, we conduct tests on the AN4 dataset—a notably small dataset—using five randomly selected seeds for both the Conformer and ConExtBiMamba models under identical hyperparameters. Our results show that ConExtBiMamba consistently outperforms the Conformer across all seeds. Specifically, the mean WER for the Conformer was 10.02, while it was significantly lower for ConExtBiMamba at just 4.54. Remarkably, ConExtBiMamba also demonstrated exceptional robustness, with a WER variance of only 1.11 compared to the Conformer’s substantially higher variance of 77.71. Observing the loss metrics, we noted that ConBiMamba is less prone to overfitting on small datasets compared with the Conformer. This suggests that ConBiMamba, like RNNs, benefits from its ability to process information recursively over time, updating only a small set of parameters with each iteration, which helps prevent overfitting in scenarios with limited data. Additionally, this framework also avoids the gradient vanishing and explosion issues typically associated with RNNs. Combining these observations with our main results, we can conclude that ConExtBiMamba effectively merges the strengths of self-attention-based models and RNN models.
| Method/Seed | 2048 | 233 | 666 | 1024 | 3407 | Average | Variance |
| Conformer | 4.0 | 8.2 | 6.1 | 4.4 | 27.4 | 10.02 | 77.71 |
| ConExtBiMamba | 3.8 | 3.8 | 4.8 | 3.8 | 6.5 | 4.54 | 1.11 |

### v6

@@ V-C Ablation Study on ConExtBiMamba
V-D Discussion
As demonstrated in Section V-B, independent Mamba or BiMamba models exhibit low performance in the ASR task, while replacing MHSA with BiMamba layers demonstrates impressive performance and outperforms the vanilla Transformer and Conformer models. Since the latter approach achieves significantly higher ASR performance by additionally employing a feed-forward network (FFN) and a residual connection compared to the independent Mamba and BiMamba models, we then explore the factors contributing to this performance improvement via ablation studies in Table XVI.
As illustrated in Section II, we consider a BiMamba layer as a weakly nonlinear module similar to MHSA [92]. We thus assume that extracting high-abstraction-level information requires greater nonlinearity than capturing low-abstraction-level information. Specifically, ASR models transcribe speech signals by understanding the context of acoustic features and aligning speech to tokens, corresponding to low-level sequential and high-level semantic information, respectively. Since a speech enhancement model focuses on low-abstraction-level spectral information, an ASR model may need higher nonlinear ability than a speech enhancement model. This assumption aligns with the results presented in Table XI, where replacing MHSA with Mamba modules in the Transformer does not consistently yield performance improvements, in contrast to their use within the Conformer model for ASR tasks. This indicates that the convolution modules in the Conformer provide more nonlinearities compared to the Transformer, enhancing the representation learning for ASR tasks.
To further validate this hypothesis, we design a visualization experiment using two synthetic datasets with different complexity levels: (1) a simple dataset containing two Gaussian clusters centered at (2,2) and (-2,-2) with a variance of 0.5; (2) a complex dataset consisting of two interleaved spirals with added Gaussian noise (). We generate 800 samples for each dataset with an 80-20 train-test split. The models are trained using Adam optimizer with a learning rate of 0.01 for 100 epochs. The decision boundaries are visualized by evaluating model predictions on a 100×100 point grid spanning the input space. In Figure 4, we use the BiMamba model to find the decision boundary for data with simple and complex distribution, respectively. We observe that BiMamba struggles to find the decision boundary for data of a complex distribution without the aid of an FFN (with ReLU activation similar to that in Transformer), which validates our assumption.
In addition, results in Table XVI indicate the effectiveness of the FFN and residual connection in a Transformer model for ASR. Similar to independent ExtBiMamba, removing the residual connection and FFN from a Transformer model leads to gradient vanishing and decreases nonlinearity, resulting in significant performance degradation for ASR. This further underpins that using BiMamba layers as a replacement for MHSA is more appropriate for speech tasks which require models to learn high-abstraction-level information than employing it independently. Recent studies proposed that training stability of Mamba models benefits significantly from the application of normalization layers [93], and suggested that the lack of inherent nonlinearity in Mamba necessitates the integration of components such as decoders or feed-forward layers to effectively handle the complexities of speech recognition [94]. These studies support our conclusion and extend it to training stability.
Similar to our investigation of Mamba in speech processing, numerous studies have explored different strategies for incorporating Mamba into ASR systems. Most of these works focus on integrating Mamba with other architectures (such as self-attention-based models) [36, 95, 96]. [35] took a different approach and demonstrated promising results with an independent Mamba encoder. Here, authors in [35] observed that the Mamba model exhibited occasional instability during training. Therefore, fine-grained hyper-parameter tuning is required with an AdamW optimizer to achieve desirable performance. Our experiments are consistent with the observation that the Mamba model was sometimes unstable during training. To ensure fair comparisons between models, we directly transferred the optimal parameters from attention-based models to the corresponding experiments involving Mamba. We attempted to replicate the experiments following the setup in [35] and found that our baseline performance improved by approximately 10% when using the AdamW optimizer. However, we were unable to exactly reproduce their results, likely due to differences in the computational environment and the random seed. Since Mamba shows consistently high performance in ASR when used as a replacement for the self-attention module rather than independently, the above underpins our claim that using Mamba to replace the self-attention module is more appropriate for tasks requiring high-abstract-level information due to greater nonlinearity compared to independent Mamba.
Based on our findings and existing literature [35, 36], we extend the scope of tasks that rely on low-abstraction-level information beyond speech enhancement to include phoneme recognition, speech separation, voice activity detection, and similar tasks. On the other hand, speech processing tasks that require high-abstraction-level information are not limited to speech recognition but also encompass speech translation, text-to-speech, and speech emotion recognition.
VI Conclusion